OpenStudy (anonymous):

Let y = 6. What is the value of 5y? A. 1 B. 11 C. 30 D. 56

OpenStudy (anonymous):

@flower2468

OpenStudy (photon336):

what do you think @pink33

OpenStudy (mathstudent55):

What does 5y mean? What operation is there between 5 and y?

OpenStudy (anonymous):

i don't know

OpenStudy (mathstudent55):

In math, a letter followed by one of more variables means multiplication. Also, variables together means multiplication.

OpenStudy (anonymous):

i thinks it b or a

OpenStudy (mathstudent55):

Examples: \(6x = 6 \times x\) \(xy = x \times y\) \(12ab = 12 \times a \times b\)

OpenStudy (mathstudent55):

Now look at your problem and the examples above. What does 5y mean?

OpenStudy (mathstudent55):

What operation is there between 5 and y?

OpenStudy (mathstudent55):

Choose the correct answer below. \(5y\) means: A. \(5 + y\) B. \(5 - y\) C. \(5 \times y\) D. \(5 \div y\)

OpenStudy (anonymous):

d

OpenStudy (mathstudent55):

5 divided by y? Did you read any of the explanations I wrote above?

OpenStudy (anonymous):

flower help me

OpenStudy (mathstudent55):

A number followed by a variable (letter) with no operation means multiplication. Variables together without an operation means multiplication. 5y means \(5 \times y\). ab means \(a \times b\) Ok so far?

OpenStudy (anonymous):

FLOWER

OpenStudy (anonymous):

thanks 4 helping me

OpenStudy (anonymous):

i guess

OpenStudy (mathstudent55):

\(5y\) means \(5 \times y\) They tell you \(\color{red}{y = 6}\), so \(5\color{red}{ y}= 5 \times \color{red}{y} = 5 \times \color{red}{6}\) What is \(5 \times 6\) ?